OpenStudy (anonymous):

Which of the following represent the zeros of f(x) = 6x3 – 31x2 + 4x + 5 ?

OpenStudy (anonymous):

options?

OpenStudy (anonymous):

f(5)=0 , so f(x)=(x-5)(6x^2-x-1), and the roots of (6x^2-x-1)=0 are x1=0.5 and x2=-1/3 Then f(x)=(x-5)(6x^2-x-1)=(x-5)6(x-0.5)(x+1/3) f(x)=(x-5)(2x-1)(3x+1)

OpenStudy (lgbasallote):

psh..who needs options

OpenStudy (lgbasallote):

look at it @Snapbacklive ...he was able to pull something out of thin air

OpenStudy (anonymous):

The cubic expression factors are: (x-5) (2 x-1) (3 x+1)

OpenStudy (lgbasallote):

on a serious note though...you should post options if you want some satisfying replies @Katiebae